OpenStudy (anonymous):

What is the factorization of the trinomial below? x2 - 10x + 25 A. (x - 25)(x + 1) B. (x - 25)(x - 1) C. (x - 5)(x - 5) D. (x - 5)(x + 5)

OpenStudy (lgbasallote):

Hint!! \[\huge x^2 - 10x + 25 \Longrightarrow x^2 - 5x - 5x + 25\]

OpenStudy (anonymous):

x2 - 10x + 25 = x^2 -10x +25 = ( x -5)( x -5)

OpenStudy (moongazer):

C.